### v1

Abstract
[…]
1 Introduction
Policy-gradient (PG) methods are a widely used class of reinforcement learning (RL) algorithms that optimize a parameterized policy directly from sampled trajectories. In LLM post-training, PG-style algorithms are now standard for improving reasoning behavior, with widely used objectives including REINFORCE Williams (1992), Proximal Policy Optimization (PPO) Schulman et al. (2017), and GRPO Shao et al. (2024).
These methods are typically implemented in an on-policy or near on-policy regime, where data generation and learning must occur sequentially after each other. This coupling becomes a major systems bottleneck for long-sequence reasoning: rollout lengths are highly variable and exhibit a long tail Hu et al. (2025)., creating pipeline bubbles and poor compute utilization in synchronous RL training.
To improve throughput, several recent systems decouple rollout generation from learning via asynchronous, pipelined training. Prior works such as asynchronous RLHF Noukhovitch et al. (2024) AReaL Fu et al. (2025), LlamaRL Wu et al. (2025), and PipelineRL Piché et al. (2025) demonstrate that overlapping generation with optimization can substantially reduce end-to-end wall-clock time. However, this efficiency comes at the cost of training quality: the sampler generates trajectories under a lagging behavior policy while the learner continues to update, making training effectively off-policy which leads to learning degradation or even collapse.
In this work, we focus on the optimization failure mode created by off-policy training and propose variance-targeted controls that can be applied on top of existing asynchronous RL frameworks. We find that the underlying mechanism behind asynchronous training instabilities is higher-variance gradient estimates. We show that the effective sample size Kish (1995) (ESS) – a well-known diagnostic for importance sampled estimates – collapses when the sampler and learner distributions significantly diverge and that this collapse predicts gradient volatility and KL spikes across reasoning tasks. Motivated by this diagnosis, we propose Variance Controlled Policy Optimization (VCPO), a simple drop-in method that directly controls the variance of the policy-gradient estimator with two complementary components: (i) an effective sample size guided step scaling that down-weights unreliable updates and (ii) a closed-form minimum-variance off-policy reward baseline based on per-trajectory gradient norms, avoiding training an auxiliary critic or value model. To make (ii) practical at scale, we present the first scalable implementation for exact per-example gradient-norm computation with negligible runtime overhead, and leverage these statistics to build a low-overhead, gradient-aware baseline for variance reduction.
[…]
2 Background
2.1 Reinforcement Learning for Large Language Models
We consider finetuning a pretrained language model policy $\pi_{\theta}$ to maximize a scalar reward $R(x,y)$ that scores a completion $y=(y_{1},\dots,y_{T})$ for a prompt $x$ (e.g., from a reward model or task-specific evaluator). Under an autoregressive policy, the likelihood of generating $y$ is
$\pi_{\theta}(y\mid x)\;=\;\prod_{t=1}^{T}\pi_{\theta}(y_{t}\mid x,y_{<t}).$
Let $\mathcal{D}$ denote a dataset (distribution) of prompts. RL optimizes the expected reward
$J(\theta)\;=\;\mathbb{E}_{x\sim\mathcal{D},\,y\sim\pi_{\theta}(\cdot\mid x)}[R(x,y)].$
On-Policy Gradient Estimator. In On-Policy RL, we estimate $\nabla_{\theta}J(\theta)$ with sequences $y$ sampled from the policy $\pi_{\theta}$ itself. This leads to the REINFORCE gradient Williams (1992)
$\nabla_{\theta}J(\theta)=\mathbb{E}_{x\sim\mathcal{D},\,y\sim\pi_{\theta}}\!\left[R(x,y)\,\nabla_{\theta}\log\pi_{\theta}(y\mid x)\right]$
$\nabla_{\theta}\log\pi_{\theta}(y\mid x)=\sum_{t=1}^{T}\nabla_{\theta}\log\pi_{\theta}(y_{t}\mid x,y_{<t})$
which motivates a surrogate loss of
$\mathcal{L}_{\text{on-policy}}(\theta)=-\mathbb{E}_{x\sim\mathcal{D},\,y\sim\pi_{\theta}(\cdot\mid x)}\!\left[R(x,y)\,\log\pi_{\theta}(y\mid x)\right]$ (1)
A standard variance reduction technique is to subtract a baseline that may depend on the state $s_{t}=(x,y_{<t})$, but not on the current action/token $y_{t}$:
$\nabla_{\theta}J(\theta)=\mathbb{E}_{x,\,y\sim\pi_{\theta}}\!\left[\sum_{t=1}^{T}A_{t}\,\nabla_{\theta}\log\pi_{\theta}(y_{t}\mid x,y_{<t})\right],$
where an advantage-like term can be written as
$A_{t}\;=\;R(s_{t},y_{t})-b(s_{t}),\qquad b(s_{t})=b(x,y_{<t}).$
In actor–critic methods such as PPO, the baseline is a learned value function $b(s_{t})=V_{\phi}(s_{t})$ and advantages $A_{t}$ are computed via GAE. Critic-free methods such as GRPO instead use structured baselines (e.g., per-prompt group normalization), avoiding the overhead of training and serving a value model. In this paper, we focus on critic-free, REINFORCE-style updates, which are widely used for LLM reasoning post-training due to their lower resource requirements.
Off-Policy Gradient Estimator. In practical LLM RL systems, rollout generation and training are inevitably mismatched, causing training to be off-policy. This mismatch can arise even without explicit asynchronous training due to numerical issues Qi et al. (2025); Xi et al. (2026), kernel and hardware issues Liu et al. (2025a), or unavoidable differences in inference and training implementations He (2025). We therefore distinguish the sampler policy $\mu$ used to generate trajectories from the current learner policy $\pi_{\theta}$ being optimized.
Importance Sampling (IS) corrects this mismatch. For a prompt $x$ and completion $y$ consisting of tokens $y_{1},\cdots,y_{T}$, define the (sequence-level) IS ratio
$w(x,y)\;\triangleq\;\frac{\pi_{\theta}(y\mid x)}{\mu(y\mid x)}=\prod_{t=1}^{T}\frac{\pi_{\theta}(y_{t}\mid y_{1}^{t-1}x)}{\mu(y_{t}\mid y_{1}^{t-1}x)}$ (2)
With importance sampling, Eq (1) can be expressed as an expectation relative to the sampler policy, which leads to the unbiased off-policy surrogate loss
$\mathcal{L}_{\text{off-policy}}(\theta)$ $=-\mathbb{E}_{x\sim\mathcal{D},\,y\sim\mu(\cdot\mid x)}\bigl[w(x,y)\,A(x,y)\,$ (3)
$\qquad\qquad\qquad\qquad\log\pi_{\theta}(y\mid x)\bigr].$
where $A(x,y)$ denotes an advantage-like signal (e.g., RbR-b or token-level advantages).
2.2 The Role of Effective Sample Size in Off-Policy RL
In practice, the core challenge of importance sampling is not bias but finite-sample variance. Known as the Curse of the Horizon Liu et al. (2018; 2020), the product structure in (2) makes $w(x,y)$ highly sensitive to small per-token probability shifts, so the resulting weights can become heavy-tailed and a few samples may dominate each update
To quantify this weight degeneracy, given per-sample gradient contributions $g_{i}$ and importance weights $w_{i}$, define standardized weights $\tilde{w}_{i}=w_{i}/\sum_{j=1}^{B}w_{j}$ and the corresponding weighted estimator $\hat{g}=\sum_{i=1}^{B}\tilde{w}_{i}g_{i}$. A common diagnostic is the effective sample size (ESS) Kong (1992); Kong et al. (1994):
$\mathrm{ESS}\;\triangleq\;\frac{\left(\sum_{i=1}^{B}w_{i}\right)^{2}}{\sum_{i=1}^{B}w_{i}^{2}}\;=\;\frac{1}{\sum_{i=1}^{B}\tilde{w}_{i}^{2}}\;\in\;[1,B].$
ESS measures how many samples effectively contribute to the weighted estimate: if weights are nearly uniform then $\mathrm{ESS}\approx B$, while if a few weights dominate then $\mathrm{ESS}\ll B$. Moreover, under mild conditions (e.g., weak correlation between $g_{i}$ and $w_{i}$), we have
$\mathrm{Var}(\hat{g})\;\approx\;\Big(\sum_{i=1}^{B}\tilde{w}_{i}^{2}\Big)\,\mathrm{Var}(g)\;=\;\frac{1}{\mathrm{ESS}}\,\mathrm{Var}(g),$
Intuitively, this means the variance of off-policy gradient estimates only matches the on-policy averaging rate with $B$ replaced by $\mathrm{ESS}$. Thus, when ESS collapses, gradient estimates become substantially higher variance, leading to high step-to-step volatility and KL. For this reason, practical RL systems and algorithms often apply clipping Schulman et al. (2017), truncation Espeholt et al. (2018), or masking which limit the influence of extreme weights, at the cost of introducing bias relative to the exact IS correction.
2.3 RL stabilization methods
While prior methods have been proposed to stabilize LLM RL in on-policy or near-on-policy settings, we discuss these works as they closely related to asynchronous training’s highly off-policy issues. These approaches broadly fall into three categories: (1) masking/clipping to suppress outlier updates, (2) algorithmic changes to the estimator/objective/baseline, and (3) system-side changes that reduce sampler–learner mismatch.
Masking / clipping mechanisms. Similar to truncation and clipping mechanisms proposed by previous works Espeholt et al. (2018); Schulman et al. (2017), recent work has propsed masking or truncating samples with outlier importance-weight ratios to prevent them from dominating updates.
Token or sequence-level Truncated Importance Sampling (TIS) constrains the IS ratio maximum. TIS has been used in large-scale asynchronous systems such as LlamaRL Wu et al. (2025) and AReal Fu et al. (2025) as well as in near on-policy RL to correct learner–sampler mismatch in modern LLM RL training stacks Yao et al. (2025b; a)
Closely related is token or sequence-level Masked Importance Sampling (MIS), which discard away updates with extreme IS ratios. IcePop proposed token-level MIS to improve stability in MoE RL Zhao et al. (2025) while other works motivate sequence-level MIS due training–inference mismatch analyses Liu et al. (2025a). Notably, DeepSeek-V3.2 uses masking sequences above a geometric-mean threshold and with negative advantages, illustrating a practical masking rule used at scale DeepSeek-AI et al. (2025).
Finally, M2PO iteratively drops token-level losses until a second-moment proxy of importance weights falls below a threshold Zheng et al. (2025c; b).
Algorithmic changes. GSPO improves training stability in MoE RL, defining importance ratios as the geometric mean of the token-level IS ratios over the sequence and performs sequence-level clipping/optimization  Zheng et al. (2025a).
Optimal Token Baseline (OTB) derives an on-policy variance-minimizing baseline and proposes a logit-based proxy (“energy”) for gradient-norms without extra backward passes Li et al. (2025).
System-side changes. A complementary direction reduces mismatch via numerical and systems alignment. Switching from BF16 to FP16 can reduce train–inference numerical inconsistencies and improve stability Qi et al. (2025), while Rollout Routing Replay (R3) stabilizes MoE RL by replaying inference-time expert routing during training Ma et al. (2025). These mitigations are orthogonal to variance-aware optimization and can be combined with our approach.
2.4 Asynchronous distributed RL
In this paper, we adopt PipelineRL-$k$ Piché et al. (2025) for our asynchronous training setup. At learner update $t$, the trainer optimizes the current policy $\pi_{\theta_{t}}$ using trajectories sampled by a (possibly stale) behavior policy $\mu=\pi_{\theta_{t^{\prime}}}$, where $k$ bounds the policy lag $t-t^{\prime}$ (with $k=0$ recovering the synchronous/on-policy regime). This induces off-policy updates whose severity increases with $k$. We also follow PipelineRL’s in-flight weight-update scheme, where sampling continues while weights are being updated. As a result, a single trajectory may span multiple policy versions.
Throughout this paper, we use sequence-level importance sampling to correct for this lag, which is the theoretically grounded importance sampling for sequence-level policy gradients, and optimize the importance-weighted surrogate objective in (3).
3 VCPO: Robust, Scalable Asynchronous Post training
3.1 Why does Asynchronous Off-Policy RL collapse?
To understand training collapse under asynchrony, we run controlled experiments with Qwen2.5-7B Base on MATH Hendrycks et al. (2021), a dataset of competition-style mathematics problems scored by exact-match final-answer correctness, and evaluate on 500 held-out problems (MATH-500). We log three signals over training: (i) the KL divergence between the rollout policy and the current learner policy (policy lag), (ii) reward / validation accuracy, and (iii) the ESS ratio, i.e., the effective sample size computed from sequence-level importance weights and normalized by the minibatch size.
Baseline choice for diagnostics. To isolate the effect of policy lag, we adopt a strong masking/clipping baseline for these didactic collapse experiments. Based on the masking and clipping sweeps in Appendix E.2, we find that sequence-level truncated importance sampling (TIS) with a higher threshold ($c=8.0$) delays collapse the longest among masking/clipping methods. We therefore use this setting as our default off-policy correction when diagnosing collapse.
Optimizer and KL regularization. Unless otherwise noted, we use AdamW with weight decay (see Appendix F.1 for details). Following prior asynchronous Fu et al. (2025); Wu et al. (2025); Espeholt et al. (2018) and synchronous Yu and others (2025) RL works, we do not incorporate explicit KL regularization loss as it constrains learning ability. We corroborate this observation with our own experiments (Appendix E.3).
Training Collapse. Figure 2 shows a representative collapse under high staleness (10 steps off-policy). As the ESS ratio collapses, updates become dominated by a few trajectories, leading to a KL explosion and an abrupt drop in training reward and validation accuracy. This supports our hypothesis that high-variance, dominated updates drives training collapse. For further evidence, see Appendix A for examples on math and reasoning tasks.
Figure 2: Sequence-Level TIS Collapse. Qwen2.5-7B Base trained on MATH task with 10-step off-policy (PipelineRL-10). The ESS ratio first degrades and then collapses, leading to a spike in rollout–policy KL divergence and a sharp drop in both training reward and validation accuracy. See Appendix F.3 for hyperparameters and training details.
3.2 Effective Sample Size Guided Step-Size Scaling for Stable Updates
If asynchronous training fails due a subset of unreliable updates, a natural idea is to reduce the impact of those updates so that optimization remains robust under policy lag. A useful analogy comes from large-batch optimization: as batch size increases, gradient noise typically decreases McCandlish et al. (2018), allowing larger stable learning rates. While SGD often follows the linear scaling rule ($\eta\propto B$Goyal et al. (2017), a widely used heuristic for adaptive optimizers such as Adam/AdamW is the square-root scaling law: increasing minibatch size by a factor $k$ permits scaling the learning rate by $\sqrt{k}$ (equivalently, $\eta\propto\sqrt{B}$Krizhevsky (2014); Hoffer et al. (2017); You et al. (2020).
In asynchronous off-policy RL, however, the nominal batch size $B$ can be misleading. As discussed in Section 2.2, importance weights may be heavy-tailed, so a batch of size $B$ can behave like a much smaller batch with only $\mathrm{ESS}$ effectively independent samples. In this regime, $\mathrm{ESS}$ is the relevant notion of effective batch size, and step-size selection should depend on $\mathrm{ESS}$ rather than on $B$.
Thus, motivated by the square-root scaling law $\eta\propto\sqrt{B}$, we propose learning-rate scaling via an effective batch-size rule
$\eta_{\text{eff}}\;\propto\;\sqrt{\rho_{\text{ess}}},\qquad\rho_{\text{ess}}\;\triangleq\;\frac{\mathrm{ESS}}{B}$ (4)
where we call $\rho_{\text{ess}}$ the effective sample size ratio.
Because even on-policy/synchronous training can have $\rho_{\text{ess}}^{\text{on}}<1$ in practice (e.g., due to inevitable training-sampling differences and finite-sample effects), we introduce an empirical on-policy reference $\rho^{{\text{on}}}_{\text{ess}}\triangleq\mathbb{E}[\rho_{\text{ess}}\mid\text{on-policy}]$ which is a constant estimated from 1 step of a on-policy run (or a running average of a few-steps) [^1] Empirically, $\rho_{\text{on}}$ is typically well-approximated by the first on-policy step.. We then rescale the step size by the relative reliability of the batch:
$\eta_{\text{eff}}\;=\;\eta\cdot\sqrt{\frac{\rho^{{\text{off}}}_{\text{ess}}}{\rho^{{\text{on}}}_{\text{ess}}}}$ (5)
This scaling preserves the on-policy step scale when $\rho_{\text{ess}}\approx\rho_{\text{on}}$, and automatically shrinks updates as $\rho_{\text{ess}}$ collapses.
Thus when training is nearly on-policy and $\mathrm{ESS}\approx B$ and $\eta_{\text{eff}}\approx\eta$; when a few samples dominate, $\mathrm{ESS}\ll B$ and $\eta_{\text{eff}}$ shrinks like $\sqrt{\mathrm{ESS}/B}$, damping the unstable, high-variance updates that coincide with KL/gradient spikes.
3.3 Minimum Variance Baselines for Off-Policy RL
Subtracting a baseline from the reward (or advantage) leaves the expected policy gradient unchanged but can substantially reduce variance. This control-variate view dates to reinforcement-comparison methods Dayan (1991) and was formalized for policy gradients Greensmith et al. (2004), with optimal baselines derived in the on-policy setting Weaver and Tao (2013).
In more recent work, systems often use simple group baselines (e.g., per-prompt mean reward) for efficiency Li et al. (2023); Shao et al. (2024), while recent work studies heuristic or approximate optimal baselines on-policy Hao et al. (2025); Li et al. (2025). In highly asynchronous training, however, importance weighting changes the variance structure. We thus derive the variance-minimizing baseline for an off-policy, importance-weighted policy-gradient estimator.
Setup. Consider the off-policy gradient estimator with a scalar baseline $b$:
$\hat{G}(b)\;=\;\frac{1}{B}\sum_{i=1}^{B}w_{i}(R_{i}-b)\,\nabla_{\theta}\log\pi_{\theta}(\tau_{i})$ (6)
where each sample $i$ contribute a score-gradient vector $g_{i}=\nabla_{\theta}\log\pi_{\theta}(\tau_{i})$, a scalar reward/advantage $R_{i}$, and an importance ratio $w_{i}$.
Minimum Variance Off-policy Baseline. As we derive in Appendix B), minimizing $\mathrm{Var}(\hat{G}(b))$ over scalar $b$ yields a closed-form solution for the off-policy optimal baseline (OPOB)
$b_{\text{OPOB}}^{\star}\;=\;\frac{\sum_{i=1}^{B}w_{i}^{2}\,\|g_{i}\|^{2}\,R_{i}}{\sum_{i=1}^{B}w_{i}^{2}\,\|g_{i}\|^{2}}.$ (7)
Compared to common group baselines (e.g., $b=\frac{1}{B}\sum_{i}R_{i}$), Eq. (7) shows that variance-optimal baselining in the off-policy regime depends on both (i) the importance weights $w_{i}$ and (ii) the gradient magnitudes $\|g_{i}\|^{2}$. Intuitively, samples that are both highly upweighted off-policy and induce large parameter changes dominate update variance and should therefore dominate the baseline.
Connections to prior baselines. Equation (7) can recover many familiar special cases. When training is on-policy ($w_{i}=1$), it reduces to the classic gradient-norm-weighted optimal baseline Greensmith et al. (2004); Weaver and Tao (2013). If, additionally, $\|g_{i}\|$ is approximately constant within a group, then $b^{\star}$ reduces to the group-mean reward baseline commonly used in practice. Our derivation shows that in the off-policy regime, both importance weights and gradient magnitudes are required for variance-optimal baselining.
3.4 Efficient Baseline-Aware Gradient Computation
A naive gradient-aware baseline requires two backward passes: one to compute per-sample gradient norms and a second to apply the resulting baseline, roughly doubling backward compute time at scale. We instead implement it with a single backward pass, adding minimal wall-clock and memory overhead.
Key observation: linearity of the baseline term. Note that expanding the equation in (6) yields
$\hat{G}(b)\;=\;\underbrace{\frac{1}{B}\sum_{i=1}^{B}w_{i}R_{i}g_{i}}_{\text{reward-weighted term}}\;-\;b\cdot\underbrace{\frac{1}{B}\sum_{i=1}^{B}w_{i}g_{i}}_{\text{score term}}.$ (8)
Thus, for any chosen $b$, the final gradient can be formed by combining two aggregated quantities:
$S_{R}\;=\;\sum_{i=1}^{B}w_{i}R_{i}g_{i},\quad S_{1}\;=\;\sum_{i=1}^{B}w_{i}g_{i}$
$\hat{G}(b)=\frac{1}{B}(S_{R}-bS_{S}).$
Score-gradient reweighting into two buffers. We take advantage of this linearity by computing per-sample score gradients $g_{i}$ once and accumulate them into two gradient buffers: (i) $G_{R}$ a buffer that accumulates the reward/advantage-weighted score gradients $R_{i}w_{i}g_{i}$, and (ii) $G_{S}$ a score buffer that accumulates score gradients $w_{i}g_{i}$. Afterwards, we form the final update by a simple linear combination of these buffers:
$G_{R}=\sum_{i}w_{i}R_{i}g_{i},\quad G_{S}=\sum_{i}w_{i}g_{i}.$
$\Rightarrow\hat{G}(b)=\frac{1}{B}\left(G_{R}-b\,G_{S}\right),\quad$
This avoids any additional backward pass to incorporate the baseline term.
Algorithm 1 Single-backward accumulation with Off-Policy Optimal Baseline baseline
1:Input: Minibatch $\{(x_{i},\tau_{i},R_{i},w_{i})\}_{i=1}^{B}$, learner policy $\pi_{\theta}$
2:Output: Baseline-aware gradient estimate $\hat{G}(b^{\star}_{\text{OPOB}})$ and scalar baseline $b^{\star}_{\text{OPOB}}$
3: Initialize gradient buffers $G_{R}\leftarrow 0$, $G_{S}\leftarrow 0$
4: Initialize baseline numerators $N\leftarrow 0$, $D\leftarrow 0$
5:for $i=1$ to $B$ do
6:  $\ell_{i}\leftarrow\log\pi_{\theta}(\tau_{i}\mid x_{i})$
7:  $g_{i}\leftarrow\nabla_{\theta}\ell_{i}$ (backprop)
8:  $s_{i}\leftarrow\|g_{i}\|_{2}^{2}$
9:  $G_{R}\leftarrow G_{R}+(w_{i}R_{i})\,g_{i}$
10:  $G_{S}\leftarrow G_{S}+w_{i}\,g_{i}$
11:  $N\leftarrow N+w_{i}^{2}\,s_{i}\,R_{i}$
12:  $D\leftarrow D+w_{i}^{2}\,s_{i}$
13:end for
[…]
16:return $\hat{G}(b^{\star}_{\text{OPOB}}),\,b^{\star}_{\text{OPOB}}$
Implementation. We implement Algorithm 1 in verl using the Megatron-LM backend so that we can efficiently obtain gradient norm statistics. As can be seen in Figure 3, our method reduces the overhead of the naive method from 100% to a 19% step time overhead while only modestly increase memory usage by 14%. See Appendix C for more implementation details.
[…]
3.5 VCPO: Variance Controlled Policy Optimization
As noted in Appendix E.2, we find that sequence-level truncated importance sampling (TIS) with a higher threshold ($c=8.0$) delays collapse the longest among masking/clipping methods. We thus combine sequence-level TIS with our effective-sample-size step-size scaling and our off-policy optimal baseline. This leads to the following surrogate loss, whose gradient equals our baseline-aware, truncated IS policy-gradient estimator
$\mathcal{L}_{\textsc{VCPO}}(\theta)$ $=-\mathbb{E}_{x\sim\mathcal{D},\,y\sim\mu(\cdot\mid x)}$ (9)
$\quad\Bigl[w^{\textsc{TIS}}(x,y)\,(R(x,y)-b_{\text{OPOB}}^{\star})\,\log\pi_{\theta}(y\mid x)\Bigr]$
$\approx-\frac{1}{B}\sum_{i=1}^{B}w^{\textsc{TIS}}_{i}\,(R_{i}-b_{\text{OPOB}}^{\star})\,\log\pi_{\theta}(y_{i}\mid x_{i}),$
where the sequence-level truncated importance weight is
$w^{\textsc{TIS}}(x,y)\;=\;\min\!\left(\mathrm{sg}\left[\frac{\pi_{\theta}(y\mid x)}{\mu(y\mid x)}\right],\;\;c\right)$
We then apply AdamW using the gradient $\nabla_{\theta}\mathcal{L}_{\textsc{VCPO}}(\theta)$, but scale learning rate $\eta$ with an ESS-scaled step size
$\eta_{\text{eff}}\;=\;\eta\cdot\sqrt{\frac{\rho^{{\text{off}}}_{\text{ess}}}{\rho^{{\text{on}}}_{\text{ess}}}}$
such each training step size scales with $\mathrm{ESS}$ accordingly. Note although we adopt TIS, we still use the unclipped IS ratios to calculate ESS.
4 Empirical Evaluation
Figure 4: GSM8K with Qwen2-1.5B under PipelineRL-12 (high policy lag). Most baselines lead to training collapse (or crash, e.g. Geometric MIS masks all sequences and has no loss), while VCPO remains stable throughout training and matches synchronous performance. Training details and hyperparameters can be found in Appendix F.2.
Figure 5: Qwen2.5-7B under PipelineRL-10 (10 Steps Off-Policy) on Countdown and MATH-500. Across both tasks, sequence level truncated importance sampling (TIS) suffers ESS-ratio collapse followed by KL/gradient instability and degraded accuracy, whereas VCPO maintains healthy ESS and stable updates, reaching synchronous performance. Training details and hyperparameters are provided in Appendix F.3.
4.1 Tasks
We evaluate VCPO across three representative RL post-training pipelines for LLM reasoning and agentic tool use. These settings require long rollouts (reasoning and multi-turn interactions with tool calls), so training is often dominated by rollout generation. As a result, strictly on-policy training underutilizes compute: the gradient updates must wait for sampling to complete, and this bottleneck worsens at scale and in sparse-reward domains that require more exploration.
We test VCPO in highly asynchronous regimes. Across all three pipelines, VCPO remains robust to substantial learner–sampler mismatch, improving throughput while preserving strong learning performance
  1. 1.
    Mathematical problem solving. We evaluate on GSM8K Cobbe et al. (2021), a benchmark of grade-school math word problems, and on MATH Hendrycks et al. (2021), a dataset of competition-style mathematics problems spanning multiple subjects and difficulty levels. For both benchmarks, we use a binary reward based on exact match of the final numeric answer and train on the official training split while evaluating on the official validation split.
  2. 2.
    General reasoning (verifiable). We construct a Countdown-style arithmetic reasoning dataset using Reasoning Gym Stojanovski et al. (2025), where rewards are verifiable by deterministic checking of the final answer. We sample 9,000 problems for training and hold out 1,000 problems for validation, following prior setups for verifiable-reward reasoning tasks Liu et al. (2025b).
  3. 3.
    Tool-integrated multi-turn reasoning. We evaluate multi-turn tool use in the SimpleTIR setting Xue et al. (2025), where the model must interleave reasoning with external tool calls. We train using the DAPO dataset Yu and others (2025) and evaluate on a held-out exam-style benchmark (AIME2025).
4.2 Metrics and Baselines
On-Policy Training as Oracle Baseline To disentangle task-learning dynamics from instabilities introduced by asynchrony, we report a fully synchronous ($k=0$) RL run as an oracle reference (matched training steps and hyperparameters).
Other Baselines. We compare against representative prior approaches spanning (i) masking/clipping-based stabilization (e.g., M2PO, geometric sequence masking, TIS), (ii) objective-level approaches such as GSPO and OTB, and (iii) numerical/systems mitigations (e.g. using fp16 precision). In this paper, we train dense models, and so MoE specific methods such as Rollout Routing Replay are not considered.
4.3 VCPO Matches Synchronous Performance with Better Throughput
Across mathematical reasoning tasks, VCPO matches the final performance of fully synchronous RL even under high asynchrony. On GSM8K and MATH-500, Figures 4 and 5 show that VCPO remains stable throughout training and reaches accuracy comparable to the synchronous reference, whereas several baselines collapse or crash in this regime. In a general reasoning task such as Countdown, Figure 5 shows that VCPO attains the same peak performance as synchronous RL with higher end-to-end throughput enabled by asynchronous training. Furthermore, component ablations show both ESS based learning rate scaling and the off-policy optimal baseline improve robustness, and combining them yields the best stability and accuracy (Appendix D).
Table 1: End-to-end training times and validation accuracy for synchronous vs. asynchronous training (lag $k$).
Method Countdown $\uparrow$ Steps GPU hours $\downarrow$
Base 1.6%
Sync (k=0k{=}0) 38.4% 400 143.2
Async (k=10k{=}10) 41.9% 400 89.6
Method MATH-500 $\uparrow$ Steps GPU hours $\downarrow$
Base 40.2%
Sync (k=0k{=}0) 72.0% 400 134.4
Async (k=10k{=}10) 71.6% 400 92.8
Method AIME 2025 $\uparrow$ Steps GPU hours $\downarrow$
Base 5.3%
Sync (k=0k{=}0) 26.7% 300 420.2
Async (k=2k{=}2) 27.8% 220 168.9
4.4 VCPO Provides State-of-the-Art Stability Under High Asynchrony
Enabling Longer Stable Training. Figure 4 summarizes our highly asynchronous training results (12 steps off-policy) on GSM8K with Qwen2-1.5B. While the synchronous baseline trains smoothly, most existing stabilization approaches fail under this level of policy lag.
In contrast, VCPO consistently avoids collapse. Across baselines: (i) ratio-based clipping (TIS) and masking (MIS) eventually collapse; (ii) algorithmic variants such as GSPO and OTB do not reliably prevent collapse; and (iii) systems-side mitigations (e.g., FP16) delay collapse but do not address the underlying instability mechanism. Although (ii) and (iii) were not designed specifically for highly asynchronous RL, we include them for completeness because they are commonly used to stabilize on-policy LLM-RL training.
See Appendix E for additional comparisons and ablations, including systematic sweeps over TIS/MIS hyperparameters, algorithmic and systems variants (GSPO, OTB, learning-rate tuning), and tasks.
Avoiding Effective Sample Size Collapse and Training Degradation. As shown in Figure 5, in both general mathematical (MATH-500) and reasoning tasks (Countdown), naively training under high asynchrony leads to ESS-ratio collapse followed by gradient instability and performance degradation. However, VCPO remains stable, maintaining a high ESS ratio, stable gradient norms, and KL levels throughout training.
4.5 VCPO Enables Fast and Stable Long-Context Multi-Turn RL
We further evaluate VCPO in a long-context, tool-integrated multi-turn setting following SimpleTIR Xue et al. (2025), where rollouts are substantially longer and training is known to be more brittle Li et al. (2025). We train with minibatch size 128 and 16 responses per prompt, using a maximum completion length of 12K tokens and up to 5 tool calls per trajectory. We train on the DAPO Yu and others (2025) dataset and evaluate on AIME 2025, see Appendix F.4 for more details.
Due to longer trajectories and prevalence of out-of-distribution tool calls, we use a lower level of asynchrony (2 steps off-policy). Figure 1 shows that unlike sequence-level TIS which collapses, VCPO improves steadily in both train reward and AIME2025 accuracy, reducing compute by $2.5\times$ compared the synchronous baseline. This demonstrates that our variance-control mechanisms transfer to long-horizon, multi-turn post-training.
5 Discussion
In this work, we introduce VCPO, a drop-in method for robust, off-policy asynchronous RL LLM post-training. We identify a consistent failure mode in highly asynchronous regimes– an effective sample size collapse that amplifies policy gradient variance–and building on this diagnosis, VCPO stabilizes training with two main contributions (i) effective-sample-size–guided step-size rescaling to automatically damp unreliable updates, and (ii) a closed-form minimum-variance baseline for importance-weighted gradients that avoids training an auxiliary value model and adds minimal overhead. Across mathematical reasoning, general reasoning, and long-context multi-turn tool use, VCPO improves robustness under substantial learner–sampler mismatch while preserving the throughput benefits of asynchronous rollouts, highlighting explicit variance control as a key ingredient for reliable asynchronous RL at scale.
Limitations and future work. We study dense Transformer policies under standard precision and a fixed set of asynchronous pipeline designs. Future work should evaluate VCPO on MoE models (where routing can amplify mismatch), under more aggressive quantization (e.g., FP8), and in substantially longer-horizon agentic settings (e.g., search/planning with sparse rewards), where additional exploration and memory mechanisms may be needed alongside variance control.
Impact Statement
More stable asynchronous RL can reduce wasted compute and enable more efficient post-training for reasoning and tool use. However, improving reliability may also increase agent capabilities and potential misuse; deployments should therefore follow domain-appropriate safety evaluations and monitoring.
Appendix A Why does the Asynchronous RL collapse?
Across model sizes and tasks, we find the same qualitative pattern: runs that remain stable maintain a healthy ESS ratio, while unstable runs exhibit a pronounced ESS-ratio collapse shortly before KL/gradient explosions and performance collapse. This supports our central diagnosis that asynchronous failure is driven by importance-weight degeneracy and dominated off-policy updates.
Figure 6: Qwen2-1.5B Base on GSM8K questions. Truncated Importance Sampling (TIS) exhibits both effective sample size (ESS) collapse and KL explosion, leading to training collapse.
Figure 7: Countdown (Qwen2-1.5B Base) under high asynchrony with TIS. Validation accuracy collapses to near zero within the first $\sim$50 training steps, indicating rapid training failure.
In addition to Figure 2, which illustrates a collapse trajectory on MATH-500, we observe the same qualitative failure pattern across another mathematical reasoning task (GSM8K) and general reasoning task (Countdown). Figure 6 shows a representative GSM8K run under high policy lag (k=12k{=}12) using truncated importance sampling (TIS): the ESS ratio degrades and then collapses (i.e., the weighted update becomes dominated by a small subset of trajectories), followed by instability in rollout–policy KL and a sharp drop in task performance. Once ESS collapses, updates become effectively single-sample, making the policy-gradient estimator extremely high variance and precipitating abrupt KL spikes. Similarly, on Countdown (Figure 7), validation accuracy collapses to near zero within the first $\sim$50 training steps, consistent with rapid weight degeneracy and dominated off-policy updates.
Appendix B Optimal Off-Policy Reward Baseline
We derive the variance-minimizing scalar reward baseline for an importance-weighted policy-gradient estimator. Throughout, we assume standard regularity conditions for the score-function gradient estimator, including (i) support: $\mu(\tau)>0$ whenever $\pi_{\theta}(\tau)>0$ so importance ratios are well-defined, and (ii) interchange of gradient and expectation so that $\mathbb{E}_{\tau\sim\pi_{\theta}}[\nabla_{\theta}\log\pi_{\theta}(\tau)]=0$. These conditions are standard in policy-gradient analyses Williams (1992); Sutton et al. (2000); Degris et al. (2012).
B.1 Single-sample derivation
Let a single trajectory sample $\tau$ have reward (or advantage) $R\in\mathbb{R}$ and score-gradient vector
$g\;\triangleq\;\nabla_{\theta}\log\pi_{\theta}(\tau).$
In the off-policy setting, trajectories are sampled from a behavior policy $\mu$, and we use the (unnormalized) importance ratio
$w\;\triangleq\;\frac{\pi_{\theta}(\tau)}{\mu(\tau)}.$
Consider the (single-sample) importance-weighted policy-gradient estimator with a constant baseline $b$:
G(b)=w(Rb)g.G(b)\;=\;w\,(R-b)\,g. (10)
Subtracting a constant baseline does not change the expected gradient because
$\mathbb{E}_{\tau\sim\mu}[wg]=\mathbb{E}_{\tau\sim\pi_{\theta}}[g]=\mathbb{E}_{\tau\sim\pi_{\theta}}[\nabla_{\theta}\log\pi_{\theta}(\tau)]=0.$
Therefore, minimizing $\mathrm{Var}(G(b))$ over $b$ is equivalent to minimizing the second moment $\mathbb{E}\|G(b)\|_{2}^{2}$.
We have
$\mathbb{E}\|G(b)\|_{2}^{2}=\mathbb{E}\!\left[w^{2}(R-b)^{2}\|g\|_{2}^{2}\right].$ (11)
Differentiating (11) with respect to $b$ yields
$\frac{\partial}{\partial b}\,\mathbb{E}\!\left[w^{2}(R-b)^{2}\|g\|_{2}^{2}\right]=-2\,\mathbb{E}\!\left[w^{2}(R-b)\|g\|_{2}^{2}\right].$
Setting the derivative to zero gives the variance-minimizing baseline
$b^{\star}=\frac{\mathbb{E}\!\left[w^{2}R\|g\|_{2}^{2}\right]}{\mathbb{E}\!\left[w^{2}\|g\|_{2}^{2}\right]}.$ (12)
B.2 Minibatch estimator and plug-in form
For a minibatch of $B$ i.i.d. samples $\{(\tau_{i},R_{i},w_{i},g_{i})\}_{i=1}^{B}$ drawn from $\mu$, the minibatch estimator is
$\hat{G}(b)\;=\;\frac{1}{B}\sum_{i=1}^{B}w_{i}(R_{i}-b)g_{i}.$
Since averaging i.i.d. samples scales variance by $1/B$, the same $b^{\star}$ in (12) minimizes $\mathrm{Var}(\hat{G}(b))$.
In practice, we use the minibatch plug-in estimate corresponding to (12):
$b^{\star}\;\approx\;\frac{\sum_{i=1}^{B}w_{i}^{2}\|g_{i}\|_{2}^{2}R_{i}}{\sum_{i=1}^{B}w_{i}^{2}\|g_{i}\|_{2}^{2}}$ (13)
Remarks.
(i) In the on-policy regime ($w\equiv 1$), Eq. (12) reduces to the classic gradient-norm-weighted optimal baseline Greensmith et al. (2004); Weaver and Tao (2013).
(ii) Relative to group-mean baselines $b=\frac{1}{B}\sum_{i}R_{i}$, Eq. (13) shows that variance-optimal baselining in the off-policy regime depends on both the importance ratios and the gradient magnitudes. In comparison, GRPO adopts the mean normalize and RLOO uses a leave-one-out estimator
$b_{\text{GRPO}}=\frac{1}{N}\sum_{i=1}^{N}R_{i},\quad b_{\text{RLOO}}^{k}=\frac{1}{N-1}\sum_{i\neq k}R_{i}$
Appendix C Efficient Baseline Aware Gradient Computations
[…]
Per-trajectory gradient-norm statistics under deferred DP synchronization.
Computing the OPOB baseline requires per-trajectory gradient norms $\|g_{i}\|_{2}$. A naive implementation would perform two backward passes: one to obtain $\|g_{i}\|_{2}$ and a second to form the baseline-adjusted policy gradient. Instead, we compute all required quantities in a single backward pass by deferring the data-parallel (DP) gradient all-reduce until after local per-trajectory statistics are computed. Concretely, we run backprop with DP communication disabled so that each worker retains its local gradients, compute the needed per-trajectory scalars locally, and then perform a single DP all-reduce on the final combined gradient.
This design preserves the FLOPs of a single backward pass while replacing the naive two-backward implementation with modest additional memory traffic (for an extra gradient buffer) and a small number of additional scalar reductions.
Two-buffer accumulation for baseline-aware gradients.
For each trajectory $\tau_{i}$, we form the per-trajectory negative log-likelihood $\mathcal{L}_{i}=-\log\pi_{\theta}(\tau_{i})$ and backpropagate once to obtain the score-gradient $g_{i}=\nabla_{\theta}\log\pi_{\theta}(\tau_{i})$. During this pass, we accumulate $g_{i}$ into two gradient buffers: (i) a reward-weighted buffer with coefficient $w_{i}R_{i}$, and (ii) a score buffer with coefficient $w_{i}$. After the loop, we compute the OPOB baseline $b^{\star}_{\mathrm{OPOB}}$ from the locally accumulated per-trajectory statistics (Algorithm 1) and form the final baseline-aware gradient $\hat{G}=\frac{1}{B}\left(G_{R}-b^{\star}_{\mathrm{OPOB}}\,G_{S}\right)$ without any additional forward/backward pass.
Appendix D Component Ablations of VCPO
We ablate VCPO into its two core variance-control components: ESS-guided step-size scaling and the off-policy optimal baseline. To quantify each component’s contribution under the same asynchronous regime and model/task, we compare: (i) TIS (sequence-level truncated IS only), (ii) TIS + ESS step-size scaling (our step-size rule applied to AdamW), (iii) TIS + OPOB (our closed-form off-policy baseline), and All runs use identical hyperparameters except for the ablated component.
Figure 8 shows that both components improve stability and performance relative to TIS alone, and are complementary.
Figure 8: Component ablations on MATH-500 under high asynchrony (Qwen2.5-7B, PipelineRL-10). We compare TIS alone, as well as adding ESS-guided step-size scaling or the off-policy optimal baseline (OPOB). Both components improve stability and accuracy.
Appendix E Further VCPO Ablation and Comparison Studies
This appendix provides additional ablations in the MATH and GSM8K setting to isolate the contributions of (i) the ESS-guided step-size rule and (ii) the optimal off-policy baseline as well compare with previous methods. Unless otherwise noted, we follow the same training setup as in the main experiments Appendix F.3).
E.1 Is Lowering the Learning Rate Sufficient for Stable Asynchronous RL?
A natural question is whether instability under high asynchrony can be addressed by simply lowering the learning rate. Our ESS-guided rule can be viewed as an automatic, data-dependent step-size schedule: it reduces the effective step size precisely when importance weights concentrate, without requiring manual learning-rate sweeps.
[…]
Figure 9: Lowering the learning rate is not a substitute for variance-aware step-size control. On MATH-500 under high asynchrony, a very small fixed learning rate (TIS + Low LR) improves stability but converges more slowly and to lower accuracy, despite exhibiting similar ESS levels. VCPO reaches higher accuracy with stable updates without learning-rate sweeps.
E.2 Are Clipping or Masking Methods Sufficient?
We next examine whether standard heuristics that directly bound importance ratios are sufficient to stabilize highly asynchronous RL. We report additional results for both masking (MIS-style) and clipping (TIS-style) applied to different ratio definitions.
Methods.
We consider three ratio variants for both masking and clipping: (i) sequence ratio (full-trajectory likelihood ratio), (ii) token ratio (per-token likelihood ratio), and (iii) geometric mean ratio (a length-normalized proxy). For masking, we zero out the update when the chosen ratio exceeds a threshold; for clipping, we cap the ratio at the threshold.
DeepSeek-v3.2 DeepSeek-AI et al. (2025) introduces a masking heuristic that is closely related to masked importance sampling, using a geometric-mean MIS with an additional sign constraint on advantages to discard unstable updates.
Thresholds.
Following prior work and practitioner guidance suggesting that bounding extreme ratios can reduce variance Yao et al. (2025b); Wu et al. (2025); Fu et al. (2025), we evaluate two representative maximum thresholds, $c\in\{2.0,8.0\}$, across all ratio variants. (We keep other hyperparameters matched to the main setup.)
In Figure 10, TIS $(a,b)$ denotes truncated importance sampling where ratios outside the interval $(a,b)$ are clipped back into $(a,b)$, while MIS $(a,b)$ denotes masked importance sampling where updates with ratios outside $(a,b)$ are discarded (i.e., assigned zero weight).
Findings and tuning sensitivity.
Figure 10 shows that these masking/clipping methods can delay collapse in some settings, but it is not a reliable substitute for VCPO under high asynchrony. In particular, sequence-level MIS is highly brittle: at both thresholds we find it can mask nearly all sequences, effectively removing the learning signal and causing training to fail. Token-level variants can be less brittle but remain unstable and sensitive to the threshold choice. Among the baselines we tested, sequence-level TIS with the larger threshold ($c=8.0$) performs best, but still underperforms VCPO and does not consistently prevent late-stage instability.
A further limitation is additional overhead of hyperparameter tuning: the importance-ratio distribution becomes broader and more heavy-tailed as model sizes increases, moving to lower preivios, so the stabilizing threshold depends strongly on the degree of off-policy-ness. In practice this requires additional sweeps over both the threshold and the ratio definition (sequence vs. token vs. geometric mean), which becomes increasingly costly when scaling to new tasks, models, or lag values.
Figure 10: Masking and clipping ablations under 12 steps off-policy on GSM8K (Qwen2-1.5B). We compare truncated importance sampling (TIS) and masked importance sampling (MIS) using sequence-level, token-level, and geometric-mean ratio statistics with thresholds $c\in\{2.0,8.0\}$. Overall, these methods bounding is highly sensitive to the threshold and fail to prevent training collapse.
M2PO-style masking.
We reproduce M2PO Zheng et al. (2025c) using the authors’ recommended masking threshold (max =0.04=0.04). Figure 11 shows that at high policy lag (PipelineRLk=12k{=}12) on GSM8K with Qwen2-1.5B, M2PO does not yield stable training: the trusted-token (mask) fraction becomes highly erratic and can collapse toward masking nearly all tokens, coinciding with KL spikes and a sharp drop in validation accuracy. We note reducing the lag to k=10k{=}10 does not reliably fix this behavior. We suspect that the commonly adopted PipelineRL style asynchronous training setup may differ from the exact regime studied in M2PO. In particular, their pipeline appears to train on exclusively stale rollouts from a fixed-lag behavior policy, whereas our system maintains a queue with mixed-staleness samples and may also include in-flight updates with partially completed rollouts. This additional heterogeneity can broaden the importance-ratio distribution and may exacerbate masking instability, so the results in Figure 11 should be interpreted as evaluating M2PO under our (more heterogeneous) PipelineRL style asynchronous stack.
Figure 11: M2PO under high asynchrony on GSM8K (Qwen2-1.5B). Top: GSM8K validation accuracy. Bottom-left: rollout–policy KL divergence (log scale). Bottom-right: trusted-token mask fraction (log scale). At k=12k{=}12 (and similarly at k=10k{=}10), the mask fraction becomes unstable and can collapse toward masking nearly all tokens, coinciding with KL instability and training collapse. Shaded regions indicate run-to-run variability.
E.3 Does KL Loss Improve Stability?
A common stabilization method is to add a KL regularizer  Schulman et al. (2017); Shao et al. (2024) toward a reference policy, either as an explicit penalty in the objective or by folding it into the reward. To test whether KL regularization alone can mitigate the collapse we observe under large policy lag, we follow prior work Shah et al. (2026) which recommends a KL penalty to the reward (i.e., no direct gradients through the KL term)
$R^{\prime}(x,y)\;=\;R(x,y)\;-\;\beta\,\mathrm{KL}\!\left(\pi_{\theta}(\cdot\mid x)\,\|\,\pi_{\mathrm{ref}}(\cdot\mid x)\right).$
We use a coefficient $\beta=0.001$ and apply this modification on top of the Seq-TIS baseline in a highly asynchronous setting ($k=12$). As shown in Figure 12, adding KL-in-reward does not prevent collapse: training still eventually becomes unstable and fails. Moreover, KL regularization reduces peak performance, yielding a lower best validation accuracy than Seq-TIS without the KL term.
Figure 12: Effect of adding a KL penalty in the reward for Seq-TIS under high asynchrony ($k=12$) on GSM8K with Qwen2-1.5B. KL-in-reward ($\beta=0.001$) does not prevent eventual collapse and also lowers the peak validation accuracy compared to Seq-TIS without KL.
E.4 Are Gradient-Norm Proxies Like Response Length or Logit “Energy” Sufficient?
Two recent works propose replacing per-trajectory gradient norms with cheaper proxies when constructing variance-reducing baselines. For example, OPO uses response length as a proxy for gradient magnitude Hao et al. (2025):
$\|g_{i}\|\propto L_{i},$ (OPO)
motivated by the intuition that longer responses contribute more log-probability terms.
OTB introduces a logit-based “energy” proxy Li et al. (2025). By defining the cumulative per-token score-function norm
$W_{t}\triangleq\sum_{j=1}^{t}\left\|\nabla_{\theta}\log\pi_{\theta}(y_{j}\mid x,y_{<j})\right\|_{2}^{2},$ (OTB)
and approximate the per-token score norm using a closed-form weight
$\hat{w}_{t}=1-2\,\pi_{\theta}(y_{t}\mid x,y_{<t})+\|\boldsymbol{\pi}_{t}\|_{2}^{2},$ (14)
where $\|\boldsymbol{\pi}_{t}\|_{2}^{2}=\sum_{v\in\mathcal{V}}\pi_{\theta}^{2}(v\mid x,y_{<t})$ is the sum of squared next-token probabilities.
While these proxies are attractive for systems reasons, not only are they are derived in the on-policy setting, but also in the on-policy setting their approximations weakly correlate with true $\|g_{i}\|$.
To disentangle the role of our ESS scaling rule, we compare OTB’s proxy with only our VCPO’s off policy optimal baseline technique. Empirically, Figure 13 shows that OTB fails to prevent learning collapse, suggesting that both length and logit “energy” are both poorly correlated with $\|g_{i}\|$, and baselines constructed from these proxies yield weaker stabilization than using the true per-trajectory gradient norms.
Further evidence can be seen in Figure 14 which shows that both OPO and OTB’s proxies do not reliably track true per-trajectory gradient norms.
Figure 13: Qwen2.5-7B Base trained on MATH-500 under high asynchrony (k=10k{=}10): the logit-energy proxy baseline (OTB) eventually becomes unstable and collapses, while our off-policy optimal baseline (OPOB) remains stable and sustains higher validation accuracy throughout training.
Figure 14: Qwen2-1.5B Base trajectories on GSM8K questions, with max response length capped to 4096 tokens
Appendix F VCPO Training Details
F.1 Optimization Hyperparameters
We list the hyperparameters for optimization used in all our experiments in Table 2.
Table 2: Optimization hyperparameters
Hyperparameter Value
Optimizer AdamW
Learning rate $1\times 10^{-6}$
Warmup steps 0
Weight decay 0.1
AdamW $\beta_{1},\beta_{2}$ 0.9, 0.999
AdamW $\epsilon$ $10^{-8}$
Gradient clipping 1.0
Entropy coefficient 0.0
KL coefficient 0.0
F.2 GSM8K
Task and reward.
We train on the GSM8K training split Cobbe et al. (2021) using a binary reward that checks whether the model’s final numeric answer exactly matches the ground truth. We evaluate on the GSM8K test set and report accuracy (exact-match) over the full set.
Asynchronous setup.
We use PipelineRL-$k$ asynchronous rollouts with maximum policy lag k=12k{=}12, and compute sequence-level importance weights between the sampler and learner policies. All methods (including our baselines) share the same sampling/training pipeline and differ only in the variance-control and/or masking rules applied to the policy-gradient update.
Table 3: Training hyperparameters for the GSM8K task
Hyperparameter Value Reference
Model Qwen2-1.5B
Learning Rate $1\times 10^{-6}$
Learning Rate Schedule Warmup Stable Decay
Learning Rate Warmup Steps 5
Learning Rate Stable Steps 395
Learning Rate Decay Steps 0
Generation Temperature 1.0
Max Prompt Token Length 1024
Response Length 2048
Number of Prompts per Batch 8
Number of Completions per Prompt 8
Batch Size (effective) 64
Number of Training Steps 400
Total Prompts Seen 3200
Total Episodes 25600
Asynchronous-specific hyperparameters
Max Staleness 12 PipelineRL-$k$ value
Number of Sampler GPUs 6
Number of Trainer GPUs 2
Max IS ratio 2.0/8.0 Max value from TIS
$\rho_{\text{ess}}^{\text{on}}$ 1.0 On-policy ESS ratio (See 5)
F.3 Countdown and MATH
Countdown (Reasoning Gym).
We build a Countdown-style arithmetic reasoning dataset using Reasoning Gym Stojanovski et al. (2025). Rewards are verifiable by deterministic checking of the final answer; we use 9,000 problems for training and 1,000 held-out problems for validation.
MATH.
We train on the official MATH training split Hendrycks et al. (2021) with an exact-match final-answer reward. We report validation accuracy on the standard MATH validation split, and additionally use the MATH-500 subset for controlled comparisons in the main paper.
Asynchronous setup.
For both tasks, we run PipelineRL-$k$ with maximum policy lag k=10k{=}10 and use sequence-level importance sampling for off-policy correction. We use $4\timesH$100 GPUs for sampling and $4\timesH$100 GPUs for training
Table 4: Training hyperparameters for the MATH and Countdown task
Hyperparameter Value Reference
Model Qwen2.5-7B
Learning Rate $1\times 10^{-6}$
Learning Rate Schedule Warmup Stable Decay
Learning Rate Warmup Steps 0
Learning Rate Stable Steps 400
Learning Rate Decay Steps 0
Generation Temperature 1.0
Max Prompt Token Length 2048
Response Length 4096
Number of Prompts per Batch 32
Number of Completions per Prompt 16
Batch Size (effective) 512
Number of Training Steps 400
Total Prompts Seen 12,800
Total Episodes 204,800
Asynchronous-specific hyperparameters
Max Staleness 10 PipelineRL-$k$ value
Number of Sampler GPUs 4
Number of Trainer GPUs 4
Max IS ratio 8.0 Max value from TIS
$\rho_{\text{ess}}^{\text{on}}$ 1.0 On-policy ESS ratio (See 5)
F.4 Multi-Turn Tool Integrated Reasoning
Task, data, and evaluation.
We follow the SimpleTIR setting Xue et al. (2025) for multi-turn tool-integrated reasoning. We train on the DAPO-17K dataset Yu and others (2025) and evaluate on held-out exam benchmarks (AIME 2024 and AIME 2025), reporting exact-match final-answer accuracy. We use Qwen2.5-7B Base (rather than newer Qwen3 variants) to reduce the risk of evaluation-set contamination.
Rollout Configuration and Asynchrony.
We generate up to 5 turns per episode with a maximum completion length of 12,288 tokens (Table F.4). We run asynchronous training with moderate policy lag (PipelineRL-2) and compute sequence-level importance weights for off-policy correction. We use $2\timesB$200 GPUs for sampling and $2\timesB$200 GPUs for training
Table 5: Training hyperparameters for the multi-turn tool integrated reasoning
Hyperparameter Value Reference
Model Qwen2.5-7B
Learning Rate $1\times 10^{-6}$
Learning Rate Schedule Warmup Stable Decay
Learning Rate Warmup Steps 0
Learning Rate Stable Steps 200
Learning Rate Decay Steps 0
Generation Temperature 1.0
Max Prompt Token Length 2048
Response Length 12,288
Number of Prompts per Batch 128
Number of Completions per Prompt 16
Batch Size (effective) 2048
Number of Training Steps 200
Total Prompts Seen 25,600
Total Episodes 409,600
Asynchronous-specific hyperparameters
Max Staleness 2 PipelineRL-$k$ value
Number of Sampler GPUs 2
Number of Trainer GPUs 2
Max IS ratio 8.0 Max value from TIS
$\rho_{\text{ess}}^{\text{on}}$ 0.55 On-policy ESS ratio (See 5)